# Secret-Key Generation from Private Identifiers under Channel Uncertainty

### v1

@@ III-B Gaussian Sources
In this subsection, we limit our discussion to a special case of setup in Section III-A where the enrollment channel is noiseless, i.e., $\tilde{X}=X$. We consider $P_{X\mathbf{Y}_{k}\mathbf{Z}_{l}}$ the joint distribution of zero-mean Gaussian random variables with a non-singular covariance matrix. Suppose that the source $X\sim\mathcal{N}(0,\sigma^{2}_{X})$, then it suffices to model the channels to the decoder and Eve as follows.
@@
$\mathbf{Y}_{k}=\mathbf{H}_{k}X+\mathbf{N}_{\mathbf{Y}_{k}},~{}\mathbf{Z}_{l}=\mathbf{H}_{l}X+\mathbf{N}_{\mathbf{Z}_{l}},$ (16)
@@
Note that the single-letter expressions characterized in Propositions 1 and 2 can be extended to the channel model (16). In proving the capacity results for the GS and CS models with Gaussian sources, we leverage Proposition 1 to show the achievability part. However, for the outer bounds, we transform the problem in (16) into a scalar Gaussian problem using sufficient statistics [39, Ch. 2]. After the transformation, it is unclear whether all constraints in Definition 3, particularly (2), remain preserved under the scalar random variables. As a result, Proposition 2 may not be directly applicable. To this end, as shown in the proof of Theorem 1, we derive new outer regions for the channel model (16) to establish the converse part of Theorem 1.
@@
To simplify the presentation of the results for Gaussian sources, we define the following rate constraints:
@@
If $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}<\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}$, then $\mathcal{R}_{G}=\mathcal{R}_{C}=\{(R_{S},R_{J},R_{L}):R_{S}=0,R_{J}\geq 0,R_{L}\geq 0\}.$
@@
See Appendix C. ∎
@@
$\mathcal{R}_{G}$ $:\displaystyle=\bigcup_{0<\alpha\leq 1}\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}$ :
@@
$\mathcal{R}_{C}$ $:\displaystyle=\bigcup_{0<\alpha\leq 1}\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}$ :
@@
The regions in (24) and (25) align with [21, Cor. 1] when [21, eq. (5)] is replaced with (6) to eliminate the quantity $I(X;Z)$. Moreover, when the storage rate is not considered, i.e., (4) is not imposed, and Eve has no side information, i.e., $\rho_{XZ}=0$), the regions in (24) and (25) simplify to [10, Th. 4.1 and 4.2].
@@ III-C Numerical Examples
We now present numerical examples for the capacity region of the GS model. Specifically, we consider three cases, with the parameters summarized as follows: (a) $\Omega_{Y}=\Omega_{Z}=1$ with $\mathbf{H}^{\intercal}_{k^{*}}=0.95$ and $\mathbf{H}^{\intercal}_{l^{*}}=0.8$, (b) $\Omega_{Y}=3$ and $\Omega_{Z}=1$ with $\mathbf{H}^{\intercal}_{k^{*}}=[0.95~{}0.95~{}0.95]$ and $\mathbf{H}^{\intercal}_{l^{*}}=0.8$, (c) $\Omega_{Y}=3$ and $\Omega_{Z}=4$ with $\mathbf{H}^{\intercal}_{k^{*}}=[0.95~{}0.95~{}0.95]$ and $\mathbf{H}^{\intercal}_{l^{*}}=[0.8~{}0.8~{}0.5~{}0.5]$. Moreover, we fix the variance of the source identifier as $\sigma^{2}_{X}=5$ for all cases.
@@
Substituting (26) into the right-hand side of (18), the optimal secret-key rate based on $R_{J}(\alpha)$ is given by $R_{S}(R_{J}(\alpha))=\frac{1}{2}\log\left(\frac{\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}(1-2^{-2{R_{J}(\alpha)}})+\sigma^{2}_{X}\mathbf{H}^{\intercal}_{l^{*}}\mathbf{H}_{l^{*}}2^{-2{R_{J}(\alpha)}}+1}{\sigma^{2}_{X}\mathbf{H}^{\intercal}_{l^{*}}\mathbf{H}_{l^{*}}+1}\right)$. Note that if $R_{J}(\alpha)\rightarrow\infty$, $R^{*}_{S}(R_{J}(\alpha))\rightarrow\frac{1}{2}\log\left(\frac{\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}+1}{\sigma^{2}_{X}\mathbf{H}^{\intercal}_{l^{*}}\mathbf{H}_{l^{*}}+1}\right)$.
@@
Figure 3 depicts the relationship of $(R_{J}(\alpha),R_{S}({R_{J}(\alpha)}))$. In this figure, Case 2 (blue curve) shows a high secret-key rate compared to the other cases. This is due to an increase in the number of antennas at the decoder, which enhances the correlation between the source and observations at the terminal. On the other hand, in Case 3 (red curve), as the number of antennas at Eve increases, the secret-key rate drops compared to Case 2 because the stronger correlation with Eve reduces the key-generation rate. Also, this case shows that even when Eve has more antennas, a positive secret-key rate is still achievable as long as $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}\geq\mathbf{H}^{\intercal}_{l^{*}}\mathbf{H}_{l^{*}}$.
IV Concluding Remarks
We studied the GS and CS models for secret-key generation from private identifiers when the authentication channels to the legitimate terminals and Eve have multiple states and outputs. In addition, the channel uncertainty for the decoder and Eve was also taken into consideration. We first derived inner and outer bounds for general discrete sources. Next, we focused on Gaussian sources, and gave full characterizations of the capacity regions in this case. Furthermore, numerical evaluations for the Gaussian case revealed how the number of antennas at the legitimate terminals and eavesdropper affects the key and storage rates.
@@ A-A Proof of Proposition 1
$R_{S}$ $~{}{\triangleq}~{}\min_{k}I(\mathbf{Y}_{k};U|V)-\max_{l}I(\mathbf{Z}_{l};U|V)-\delta,$
$R_{J}$ $~{}{\triangleq}~{}\max_{k}I(\tilde{\mathbf{X}};U|V,\mathbf{Y}_{k})+\max_{k}I(\tilde{\mathbf{X}};V|\mathbf{Y}_{k})+5\delta,$
@@
$~{}~{}~{}-I(\tilde{\mathbf{X}};U|X)+\min_{k}I(\mathbf{Y}_{k};V)-\min_{l}I(\mathbf{Z}_{l};V)+4\delta.$
@@
$R_{v}$ ${\triangleq}~{}I(\tilde{\mathbf{X}};V)+\delta,~{}~{}{R_{J_{v_{1}}}}{\triangleq}\max_{k}I(\tilde{\mathbf{X}};V|\mathbf{Y}_{k})+2\delta,$
$R_{u}$ ${\triangleq}~{}I(\tilde{\mathbf{X}};U|V)+\delta,~{}~{}{R_{J_{u_{1}}}}{\triangleq}\max_{k}I(\tilde{\mathbf{X}};U|\mathbf{Y}_{k},V)+3\delta,$
@@
$\leq\mathbb{P}\{\mathcal{E}_{1}\}+\mathbb{P}\{\mathcal{E}_{2}\}+\mathbb{P}\{\mathcal{E}_{3}\cap(\mathcal{E}_{1}\cup\mathcal{E}_{2})^{c}\}+\mathbb{P}\{\mathcal{E}_{4}\}+\mathbb{P}\{\mathcal{E}_{5}\}.$
The first and second terms vanish by the covering lemma [43, Lemma 3.3] since $R_{v}>I(\tilde{\mathbf{X}};V)$ and $R_{u}>I(\tilde{\mathbf{X}};U|V)$, respectively. The third term vanishes by Markov lemma [39, Lemma 15.8.1]. The last two terms vanish by the packing lemma [43, Lemma 3.1], since the rate of index $\hat{j}_{v_{2}}$ is less than $\min_{k}I(\mathbf{Y}_{k};V)$ and that of index pair $(\hat{j}_{u_{2}},\hat{j}_{u_{3}})$ is less than $\min_{k}I(\mathbf{Y}_{k};U|V)$, respectively. Hence, we have
@@
Before we analyze the constraints (3), (4), (5), and (6) in Definition 3, we state the following lemmas.
@@
(Adapted from [34, Lemma A.1]) If the inequality $I(\mathbf{Z}_{l};U|V)<I(\mathbf{Z}_{l^{\prime}};U|V)$ holds for $l,l^{\prime}\in\mathcal{L}$, then there exists a random variable $A$ such that the equality $I(\mathbf{Z}_{l},A;U|V)=I(\mathbf{Z}_{l^{\prime}};U|V)$ and the Markov chain $V-U-\tilde{\mathbf{X}}-(\mathbf{Z}_{l},\mathbf{Z}_{l^{\prime}})-A$ are satisfied.
An Intermediate Step: For any $l\in\mathcal{L}$, by Lemma 2, there exists $A$ such that for
$\tilde{\mathbf{Z}}\triangleq(Z_{l},A),$ (28)
@@
Moreover, define a binary random variable $T$, which takes $1$ if $(U^{n},\tilde{\mathbf{X}}^{n},\tilde{\mathbf{Z}}^{n})\in\mathcal{T}^{n}_{\delta}$ and 00 otherwise. For large enough $n$, it holds that
@@
with $\tilde{\delta}_{n}\downarrow 0$ as $\delta\downarrow 0$ and $n\rightarrow\infty$. This follows because the pair $(U^{n},\tilde{\mathbf{X}}^{n})$ is jointly typical with probability approaching one, as shown in (27), and $\tilde{\mathbf{Z}}^{n}$ is i.i.d. generated according to $\prod_{t=1}^{n}P_{\tilde{\mathbf{Z}}_{t}|\tilde{\mathbf{X}}_{t}}$ from (29), and thus (30) follows by applying the Markov lemma [39, Lemma 15.8.1]. Similarly, we have joint typicality of $(V^{n},\tilde{\mathbf{Z}}^{n})$ as $(V^{n},\tilde{\mathbf{X}}^{n})$ is jointly typical with high probability. These properties are applied in proving the next lemma, which plays a key role in the analyses of the secret-key uniformity and secrecy-leakage.
@@
$\geq H(J_{u_{2}}|J_{u_{1}},J_{v_{1}},J_{v_{2}},\mathbf{Z}^{n}_{l},A^{n},\mathcal{C}_{n})$
=$\overset{\rm(a)}{=}H(J_{u_{2}}|J_{u_{1}},J_{v_{1}},J_{v_{2}},\tilde{\mathbf{Z}}^{n},\mathcal{C}_{n})$
@@
=$\overset{\rm(h)}{=}n(I(\tilde{\mathbf{X}};U)-I(\tilde{\mathbf{Z}};U|V))$
@@
=$\overset{\rm(i)}{=}n(\min_{k}I(\mathbf{Y}_{k};U|V)-\max_{l}I(\mathbf{Z}_{l};U|V)-\delta-\xi_{n})$
@@
$=2^{-n(I(\tilde{\mathbf{X}};U)+H(\tilde{\mathbf{Z}}|U)-2\epsilon_{\delta})},$
(f) follows because, as shown in the intermediate step, $(V^{n},\tilde{\mathbf{Z}}^{n})$ is jointly typical with high probability and thus $H(\tilde{\mathbf{Z}}^{n}|V^{n},\mathcal{C}_{n})\leq n(H(\tilde{\mathbf{Z}}|V)+\gamma_{n})$ (cf. [44, eq. (16)]) and $\delta^{\prime}_{n}\triangleq\tilde{\delta}_{n}(I(\tilde{\mathbf{X}};U)+H(\tilde{\mathbf{Z}}|U))+2(1-\tilde{\delta}_{n})\epsilon_{\delta}+\delta_{n}$, (g) is due to the Markov chain (29) and $H(J_{u_{1}}|\mathcal{C}_{n})\leq n{R_{J_{u_{1}}}}$, $H(J_{v_{1}}|\mathcal{C}_{n})\leq~{}n{R_{J_{v_{1}}}}$, $H(J_{v_{2}}|\mathcal{C}_{n})\leq n(R_{v}-R_{J_{v_{1}}})=n(\min_{k}I(\mathbf{Y}_{k};V)-\delta)$, (h) is due to the Markov chain $V-U-\tilde{\mathbf{X}}-\mathbf{Y}_{k}$, (i) follows from $I(\tilde{\mathbf{Z}};U|V)=\max_{l}I(\mathbf{Z}_{l};U|V)$ and $\xi_{n}~{}\triangleq~{}3\delta+\gamma_{n}+{\delta^{\prime}_{n}}$. ∎

### v2

@@ III-B Gaussian Sources
In this subsection, we limit our discussion to a special case of setup in Section III-A where the enrollment channel is noiseless, i.e., $\tilde{\mathbf{X}}=X$. We consider $P_{X\mathbf{Y}_{k}\mathbf{Z}_{l}}$ the joint distribution of zero-mean Gaussian random variables with a non-singular covariance matrix. Suppose that the source $X\sim\mathcal{N}(0,\sigma^{2}_{X})$, then it suffices to model the channels to the decoder and Eve as follows.
@@
$\mathbf{Y}_{k}=\mathbf{H}_{k}X+\mathbf{N}_{\mathbf{Y}_{k}},~\mathbf{Z}_{l}=\tilde{\mathbf{H}}_{l}X+\mathbf{N}_{\mathbf{Z}_{l}},$ (17)
@@
Remark 2.
In the case where the enrollment channel is noisy, i.e., $\tilde{\mathbf{X}}\neq X$, the noise covariance matrices of the involved channels are not positive definite in general. This prevents the use of Cholesky decomposition to normalize them to identity matrices, and thus the channel models described in Lemma 1 may no longer be applicable.
Note that the single-letter expressions characterized in Propositions 1 and 2 can be extended to the channel model (17). To derive a closed-form analytical (parametric) expression for Gaussian sources, we directly leverage Proposition 1 to show the achievability. In the converse, we transform the problem in (17) into a scalar Gaussian problem using sufficient statistics [46, Ch. 2], which helps avoid the complexity of working with vector random variables. However, after the transformation, it is unclear whether all constraints in Definition 3, particularly (2), remain preserved under the scalar random variables. As a result, Proposition 2 may not hold when the vector random variables are replaced with scalar ones. To this end, as shown in the proof of Theorem 1, we derive new outer regions for the channel model (17) using scalar variables to establish the converse part of Theorem 1.
@@
To simplify the presentation of the results for Gaussian sources, we define the following rate constraints, where $\alpha\in(0,1]$ serves as a tuning parameter that adjusts the variance of the auxiliary Gaussian random variable. For further details, the reader is referred to (60).
@@
If $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}<\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}$, then
$\mathcal{R}_{\rm G}=\mathcal{R}_{\rm C}=\{(R_{S},R_{J},R_{L}):R_{S}=0,R_{J}\geq 0,R_{L}\geq 0\}.$ (25)
@@
The proof is provided in Appendix C and includes the achievability and converse parts. For the achievability, we set the test channel $P_{U|X}$ to be an AWGN channel and then apply Weinstein–Aronszajn Identity [52, Appx. B] to calculate the mutual information with vector random variables. Finally, we use Lemma 5 to show that the optimal inner region is achieved when the indices of the channels to the decoder and Eve are $k^{*}$ and $l^{*}$, respectively. For the converse, we begin by invoking the sufficient statistics [46] to convert vector variables to scalar ones. Next, we derive a single-letter characterization of the outer bound using these scalar variables, which is then used to determine the parametric expressions for the Gaussian case. The proof employs a technique based on Fisher information, introduced in [49]. In the final step, we again apply Lemma 5 to derive the outer region valid for an arbitrary pair $(k,l)$, which is obtained when the decoder and Eve observe the channels indexed by $k^{*}$ and $l^{*}$ as well, coinciding with the optimal inner bound. ∎
In Theorem 1, the condition $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}\geq\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}$ indicates that the channel power gain of the worst link to the decoder is at least as large as that of the best link to Eve. In the single-antenna case, i.e., $|\mathcal{K}|=1=|\mathcal{L}|$, this condition corresponds to physically degraded channels, where the channel to Eve is physically degraded with respect to the channel to the decoder.
Unlike the discrete sources, the inner and outer bounds for the Gaussian sources coincide. This is because, in the outer bound, the variable involved in the optimization is a scalar parameter, and rate constraints are given by logarithmic functions of the optimization parameter, $\alpha$, and the values of channel power gains $\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}$ and $\tilde{\mathbf{H}}^{\intercal}_{l}\tilde{\mathbf{H}}_{l}$. These functions are monotonic with respect to the channel power gains for an arbitrary $\alpha$. As a result, the order of intersection and union does not matter and can be swapped, which enable us to take the intersection over channel states $(k,l)$ for each rate constraint and determine the saddle point $(k^{*},l^{*})$ at which the outer bound matches the inner bound.
@@
$\mathcal{R}_{\rm G}$ $=\bigcup_{0<\alpha\leq 1}\Big\{(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:$
@@
$\mathcal{R}_{\rm C}$ $=\bigcup_{0<\alpha\leq 1}\Big\{(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:$
@@
The regions in (26) and (27) align with [22, Cor. 1] when [22, eq. (5)] is replaced with (6) to eliminate the quantity $I(X;Z)$. Moreover, when the storage rate is not considered, i.e., (4) is not imposed, and Eve has no side information, i.e., $\rho_{XZ}=0$, the regions in (26) and (27) simplify to [11, Th. 4.1 and 4.2].
@@ III-C Numerical Examples
We begin by presenting numerical calculations that illustrate the relationship between the secret-key and storage rates in the GS model, and then proceed to compare the secret-key and privacy-leakage rates of the GS and CS models under the same storage rate.
For investigating the relation of the secret-key and storage rates, we consider three cases, with the parameters summarized as follows: 1. $\Omega_{Y}=\Omega_{Z}=1$ with $\mathbf{H}^{\intercal}_{k^{*}}=0.95$ and $\tilde{\mathbf{H}}^{\intercal}_{l^{*}}=0.8$, 2. $\Omega_{Y}=3$ and $\Omega_{Z}=1$ with $\mathbf{H}^{\intercal}_{k^{*}}=[0.95~0.95~0.95]$ and $\tilde{\mathbf{H}}^{\intercal}_{l^{*}}=0.8$, and 3. $\Omega_{Y}=3$ and $\Omega_{Z}=4$ with $\mathbf{H}^{\intercal}_{k^{*}}=[0.95~0.95~0.95]$ and $\tilde{\mathbf{H}}^{\intercal}_{l^{*}}=[0.8~0.8~0.5~0.5]$. Moreover, we fix the variance of the source identifier as $\sigma^{2}_{X}=5$ for all cases.
@@
Substituting (28) into the right-hand side of (19), the optimal secret-key rate based on $R_{J}(\alpha)$ is given by $R_{S}(R_{J}(\alpha))=\frac{1}{2}\log\left(\frac{\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}(1-2^{-2{R_{J}(\alpha)}})+\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}2^{-2{R_{J}(\alpha)}}+1}{\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}+1}\right)$. Note that if $R_{J}(\alpha)\rightarrow\infty$, $R^{*}_{S}(R_{J}(\alpha))\rightarrow\frac{1}{2}\log\left(\frac{\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}+1}{\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}+1}\right)$.
@@
Figure 3(a) depicts the relation of $(R_{J}(\alpha),R_{S}({R_{J}(\alpha)}))$. In this figure, Case 2 (blue curve) shows a high secret-key rate compared to the other cases. This is due to an increase in the number of antennas at the decoder, which enhances the correlation between the source and observations at the terminal. On the other hand, in Case 3 (red curve), as the number of antennas at Eve increases, the secret-key rate drops compared to Case 2 because the stronger correlation with Eve reduces the key-generation rate. Also, Case 3 shows that even when Eve has more antennas, a positive secret-key rate is still achievable as long as $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}\geq\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}$.
Figure 3(b) presents the secret-key and storage rates for a given $\alpha$, focusing on Case 3, where the maximum secret-key rate reaches 0.2771 (cf. Fig. 3(a)). As $\alpha\to 0$, the storage rate grows unbounded, reflecting the absence of encoding, while the secret-key rate is maximized. In contrast, as $\alpha\to 1$, both rates approach zero. According to (60), this is because $U$ is highly correlated with $X$ when $\alpha\to 0$, leading to a high secret-key rate, whereas $U$ becomes independent of $X$ when $\alpha=1$, resulting in zero secret-key rate.
Figures 3(c) and 3(d) respectively compare the secret-key and privacy-leakage rates between the GS and CS models for Case 2, under the same values of storage rates. In the low storage rate regime, the GS model results in a higher secret-key rate than the CS model, but at the cost of greater privacy leakage, highlighting a trade-off between these two security metrics. In the high storage rate regime, both models achieve the same secret-key rate, but the GS model still incurs greater privacy leakage than the CS model with the difference equal to the secret-key rate. This occurs because, in the CS model, the concealed data (with rate equal to the secret-key rate) reveals no information about the source identifier. These results suggest that in practical system designs, where the storage space is fixed, the GS model may be preferred in the low storage rate regime when maximizing the secret-key rate is important, while the CS model is preferable for minimizing privacy leakage. In the high storage rate regime, the CS model becomes the preferred option as it can achieve the same secret-key rate as the GS model but with lower privacy leakage.
Figure 3: (a) the relation of storage and secret-key rates in the GS model, (b) the secret-key and storage rates for a given value of $\alpha$ in the GS model, and for given a storage rate, a comparison of the secret-key and privacy-leakage rates in the GS and CS models are shown in (c) and (d), respectively.
IV Concluding Remarks and Future Directions
We studied secret-key generation from private identifiers under channel uncertainty and multiple-output settings. This setup addresses authentication robustness against eavesdroppers in scenarios where the legitimate terminals lack exact CSI and Eve may use multiple antennas to improve signal reception. We derived inner and outer bounds for discrete sources and provided a full capacity characterization for Gaussian sources. The main technical contributions lie in proving the inner bound for discrete memoryless sources and the outer bound for the Gaussian case.
To prove the inner bound for discrete sources, we first extend the technique used in [40] for compound wiretap channels to ensure that the generated secret key is uniform and remains secret from Eve’s observation for any channel state. For the outer bound in the Gaussian case, we first employ sufficient statistics to convert the vector problem into a scalar one, so that we can use the degraded property of the scalar Gaussian random variables to derive a single-letter characterization of the outer region. Then, we apply the single-letter characterization to derive the parametric expression for the Gaussian case with Fisher information-based techniques playing a crucial role in the derivation.
We also performed numerical evaluations for the Gaussian case to illustrate how changes in the number of antennas at the legitimate terminals and the eavesdropper affect the trade-offs between secret-key and storage rates, and to compare the secret-key and privacy-leakage rates of the GS and CS models under the same storage rate. The first set of results indicates that increasing the number of antennas at the decoder leads to a higher secret-key rate, while adding antennas at Eve reduces the secret-key rate. Nevertheless, even if Eve has more antennas, a positive secret-key rate remains achievable as long as the worst-case channel power gain at the decoder exceeds the best-case channel power gain at Eve. The second set of results shows that in the low storage-rate regime, the GS model achieves a higher secret-key rate, whereas the CS model offers better privacy-leakage performance. In contrast, in the high storage-rate regime, the CS model proves to be the more favorable choice as it provides the same secret-key rate as the GS model with lower privacy leakage.
A natural extension of this work is to characterize the capacity region for Gaussian sources under noisy enrollment channels. As noted in Lemma 2, since we may not be able to model the covariance matrices of the independent noises as identity matrices, the analysis will become more involved compared to that of Theorem 1. This arises because the scalar problem obtained by transforming the original vector problem using sufficient statistics results in more complicated forms than the expressions in Lemma 7. Extending the scenario to the case of vector Gaussian sources is also an interesting topic. Another possible avenue is to include user identification as studied in [23, 24, 25] and see how the identification rate influences the capacity region.
@@ A-A Proof of Proposition 1
$R_{S}$ $~\triangleq~\min_{k}I(\mathbf{Y}_{k};U|V)-\max_{l}I(\mathbf{Z}_{l};U|V)-\delta,$ (29)
$R_{J}$ $~{\triangleq}~\max_{k}I(\tilde{\mathbf{X}};U|V,\mathbf{Y}_{k})+\max_{k}I(\tilde{\mathbf{X}};V|\mathbf{Y}_{k})+5\delta,$ (30)
@@
$~~~-I(\tilde{\mathbf{X}};U|X)+\min_{k}I(\mathbf{Y}_{k};V)-\min_{l}I(\mathbf{Z}_{l};V)+4\delta.$ (31)
@@
$R_{v}$ ${\triangleq}~I(\tilde{\mathbf{X}};V)+\delta,~~{R_{J_{v_{1}}}}{\triangleq}\max_{k}I(\tilde{\mathbf{X}};V|\mathbf{Y}_{k})+2\delta,$ (32)
$R_{u}$ ${\triangleq}~I(\tilde{\mathbf{X}};U|V)+\delta,~~{R_{J_{u_{1}}}}{\triangleq}\max_{k}I(\tilde{\mathbf{X}};U|\mathbf{Y}_{k},V)+3\delta,$ (33)
@@
$\leq\mathbb{P}\{\mathcal{E}_{1}\}+\mathbb{P}\{\mathcal{E}_{2}\}+\mathbb{P}\{\mathcal{E}_{3}\cap(\mathcal{E}_{1}\cup\mathcal{E}_{2})^{c}\}+\mathbb{P}\{\mathcal{E}_{4}\}+\mathbb{P}\{\mathcal{E}_{5}\}.$ (34)
The first and second terms vanish by the covering lemma [51, Lemma 3.3] since $R_{v}>I(\tilde{\mathbf{X}};V)$ and $R_{u}>I(\tilde{\mathbf{X}};U|V)$, respectively. The third term vanishes by Markov lemma [46, Lemma 15.8.1]. The last two terms vanish by the packing lemma [51, Lemma 3.1], since the rate of index $\hat{j}_{v_{2}}$ is less than $\min_{k}I(\mathbf{Y}_{k};V)$ and that of index pair $(\hat{j}_{u_{2}},\hat{j}_{u_{3}})$ is less than $\min_{k}I(\mathbf{Y}_{k};U|V)$, respectively. Hence, we have
@@
Before we analyze the constraints (3), (4), (5), and (6) in Definition 3, we state two lemmas. The first one, Lemma 2, is an extended version of [40, Lemma A.1] to incorporate conditional mutual information.
@@
If the inequality $I(\mathbf{Z}_{l};U|V)<I(\mathbf{Z}_{l^{\prime}};U|V)$ holds for $l,l^{\prime}\in\mathcal{L}$, then there exists a vector random variable $\mathbf{A}$ such that the equality $I(\mathbf{Z}_{l},\mathbf{A};U|V)=I(\mathbf{Z}_{l^{\prime}};U|V)$ and the Markov chain $V-U-\tilde{\mathbf{X}}-(\mathbf{Z}_{l},\mathbf{Z}_{l^{\prime}})-\mathbf{A}$ are satisfied.
Proof.
Let $B$ be a binary random variable taking values $l$ and $l^{\prime}$ with probabilities $p$ and 1p1-p, respectively, where $0\leq p\leq 1$, and assume that $B$ is independent of all other random variables. Define $\mathbf{A}\triangleq(\mathbf{Z}_{B},B)$ and $\Gamma(p)=I(\mathbf{Z}_{l},\mathbf{A};U|V)$. Due to the independence of $B$, we have
$\Gamma(p)=pI(\mathbf{Z}_{l};U|V)+(1-p)I(\mathbf{Z}_{l},\mathbf{Z}_{l^{\prime}};U|V).$ (36)
Now observe that $\Gamma(1)<I(\mathbf{Z}_{l^{\prime}};U|V)\leq\Gamma(0)$, where the first inequality follows by the assumption $I(\mathbf{Z}_{l};U|V)<I(\mathbf{Z}_{l^{\prime}};U|V)$. Due to the continuity of the function $\Gamma(p)$ for all $p\in[0,1]$, there exists a $p^{*}\in[0,1]$ such that $\Gamma(p^{*})=I(\mathbf{Z}_{l^{\prime}};U|V)$, and thus the equality $I(\mathbf{Z}_{l},\mathbf{A};U|V)=I(\mathbf{Z}_{l^{\prime}};U|V)$ is satisfied with $\mathbf{A}=(\mathbf{Z}_{B^{*}},B^{*})$ and $B^{*}$ taking the values $l$ and $l^{\prime}$ with probabilities $p^{*}$ and $1-p^{*}$, respectively. Also, this choice of $\mathbf{A}$ ensures that the Markov chain $V-U-\tilde{\mathbf{X}}-(\mathbf{Z}_{l},\mathbf{Z}_{l^{\prime}})-\mathbf{A}$ is satisfied. ∎
[…]
Intermediate Step: For any $l\in\mathcal{L}$, by Lemma 2, there exists $\mathbf{A}$ such that for
$\tilde{\mathbf{Z}}\triangleq(\mathbf{Z}_{l},\mathbf{A}),$ (37)
@@
Moreover, define a binary random variable $T$, which takes $1$ if $(U^{n},\tilde{\mathbf{X}}^{n},\tilde{\mathbf{Z}}^{n})\in\mathcal{T}^{n}_{\delta}$ and 0 otherwise. For large enough $n$, it holds that
@@
with $\tilde{\delta}_{n}\downarrow 0$ as $\delta\downarrow 0$ and $n\rightarrow\infty$. This follows because the pair $(U^{n},\tilde{\mathbf{X}}^{n})$ is jointly typical with probability approaching one, as shown in (35), and $\tilde{\mathbf{Z}}^{n}$ is i.i.d. generated according to $\prod_{t=1}^{n}P_{\tilde{\mathbf{Z}}_{t}|\tilde{\mathbf{X}}_{t}}$ from (38), and thus (39) follows by applying the Markov lemma [46, Lemma 15.8.1]. Similarly, we have joint typicality of $(V^{n},\tilde{\mathbf{Z}}^{n})$ as $(V^{n},\tilde{\mathbf{X}}^{n})$ is jointly typical with high probability. These properties are applied in proving the next lemma, which plays a key role in the analyses of the secret-key uniformity and secrecy-leakage.
@@
$\geq H(J_{u_{2}}|J_{u_{1}},J_{v_{1}},J_{v_{2}},\mathbf{Z}^{n}_{l},\mathbf{A}^{n},\mathcal{C}_{n})$
$\overset{\rm(a)}{=}H(J_{u_{2}}|J_{u_{1}},J_{v_{1}},J_{v_{2}},\tilde{\mathbf{Z}}^{n},\mathcal{C}_{n})$
@@
$\overset{\rm(h)}{=}n(I(\tilde{\mathbf{X}};U)-I(\tilde{\mathbf{Z}};U|V))$
@@
$\overset{\rm(i)}{=}n(\min_{k}I(\mathbf{Y}_{k};U|V)-\max_{l}I(\mathbf{Z}_{l};U|V)-\delta-\xi_{n})$
@@
$=2^{-n(I(\tilde{\mathbf{X}};U)+H(\tilde{\mathbf{Z}}|U)-2\epsilon_{\delta})},$ (42)
(f) follows because, as shown in the intermediate step, $(V^{n},\tilde{\mathbf{Z}}^{n})$ is jointly typical with high probability and thus $H(\tilde{\mathbf{Z}}^{n}|V^{n},\mathcal{C}_{n})\leq n(H(\tilde{\mathbf{Z}}|V)+\gamma_{n})$ (cf. [53, eq. (16)]) and $\delta^{\prime}_{n}\triangleq\tilde{\delta}_{n}(I(\tilde{\mathbf{X}};U)+H(\tilde{\mathbf{Z}}|U))+2(1-\tilde{\delta}_{n})\epsilon_{\delta}+\delta_{n}$, (g) is due to the Markov chain (38) and $H(J_{u_{1}}|\mathcal{C}_{n})\leq n{R_{J_{u_{1}}}}$, $H(J_{v_{1}}|\mathcal{C}_{n})\leq~n{R_{J_{v_{1}}}}$, $H(J_{v_{2}}|\mathcal{C}_{n})\leq n(R_{v}-R_{J_{v_{1}}})=n(\min_{k}I(\mathbf{Y}_{k};V)-\delta)$, (h) is due to the Markov chain $V-U-\tilde{\mathbf{X}}-\mathbf{Y}_{k}$, (i) follows from $I(\tilde{\mathbf{Z}};U|V)=\max_{l}I(\mathbf{Z}_{l};U|V)$ and $\xi_{n}~\triangleq~3\delta+\gamma_{n}+{\delta^{\prime}_{n}}$. ∎